Of 150 people, 60 are men and 90 are women, 30 of the men and 60 of the women
read news paper daily. A person is selected at random from the lot. What is the chance that selected would be a man or read newspaper?

Answer by Edwin McCravy(20067)

P(Man) = 60/150 = 2/5 = 0.4
P(Man AND Reads Newspaper) = 30/150 = 1/5 = 0.2 
P(Read Newspaper) = (30+60)/150 = 90/150 = 3/5 = 0.6

P(Man OR Reads Newspaper) = P(Man) + P(Reads Newsaper) - P(Man AND Reads Newspaper)
P(Man OR Reads Newspaper) = 0.4 + 0.6 - 0.2 = 0.8

Edwin